# PinRec: Outcome-Conditioned, Multi-Token Generative Retrieval for Industry-Scale Recommendation Systems

### v2

@@ Abstract.
This paper introduces PinRec, a novel generative retrieval model developed for applications at Pinterest. PinRec utilizes outcome-conditioned generation, enabling modelers to specify how to balance various outcome metrics, such as the number of saves and clicks, to effectively align with business goals and user exploration. Additionally, PinRec incorporates multi-token generation to enhance output diversity while optimizing generation. Our experiments demonstrate that PinRec can successfully balance performance, diversity, and efficiency, delivering a significant positive impact to users using generative models. This paper marks a significant milestone in generative retrieval, as it presents, to our knowledge, the first rigorous study on implementing generative retrieval at the scale of Pinterest.
@@ 1. Introduction
Pinterest is a visual discovery platform where over 550 million monthly active users search for, save, and shop for ideas, represented by Pins (Pinterest, 2024a). Given the massive volume of Pins — for instance, users save over 1.5 billion Pins per week (Pinterest, 2024b) — it is essential to provide personalized recommendations in Pinterest feeds and search results. To achieve this, recommender systems often leverage a two-stage approach: retrieval and ranking. Retrieval models select thousands of potentially relevant candidate items, then ranking models prioritize the top items to display to the user. Since ranking models depend on the retrieval results, it is important for retrieval models to identify diverse, fresh, and relevant items.
@@
Traditionally, retrieval methods employ “two-tower” models. The query tower computes query embeddings, while the item tower computes item embeddings. These models retrieve items with the highest similarity scores given a query embedding. As an alternative to two-tower models, recent work has proposed generative retrieval. Generative retrieval models leverage a sequence of tokens using generative models, such as transformers (Vaswani et al., 2017), and then use those tokens to retrieve items. Generative retrieval has shown promising results in academic settings given its powerful ability to understand user activity history and generate candidates that existing two-tower models may not be able to retrieve (Kang and McAuley, 2018a; Li et al., 2023; Yang et al., 2024). Nevertheless, to our knowledge, generative retrieval has not been attempted as successfully for large-scale industrial applications yet.
@@
Using generative retrieval models for large scale recommender systems poses many challenges at the scale of Pinterest. Given their increased complexity and sequential generation, generating a set of recommended items requires higher computational cost than computing an embedding with two-tower models. Further, the generative model should be capable of efficiently optimizing for different engagement outcomes because (a) business needs are constantly evolving to optimize combinations of different metrics and (b) as importantly, we wish for users to engage with content in the most optimal way for their growth on the platform. For instance, we would want to approach recommendations for a user with a propensity for low-effort engagement differently than for core users; through which recommendations could we enable the maximum progress in their user journey on Pinterest? To answer this question requires a malleable technique that can adapt to different circumstances, goals, and demands. Finally, the generated item sequences must possess sufficient diversity to facilitate users’ exploration of varied ideas; one-step generation of a single item representation may not suffice.
@@
To our knowledge, there has been no large scale generative retrieval model that has addressed all of the aforementioned challenges. Most methods developed within academic settings (Kang and McAuley, 2018a; Hou et al., 2022; Li et al., 2023) have not proven effective in meeting the latency and cost requirements necessary to handle web-scale traffic. While there are a few recent studies to apply generative retrieval to production, they focus on optimizing one objective (Zhai et al., 2024) and do not demonstrate optimal optimization for different target outcomes.
@@
In this paper, we present PinRec, a generative retrieval technique for real-world recommender systems. Unlike previous methods, PinRec addresses and resolves three crucial criteria for generation of optimal recommendations: multifaceted generative capabilities, computational efficiency, and item diversity. Our main contributions in the paper are summarized below.
  • Outcome-Conditioned Generative Retrieval: Given evolving business needs and to allow for different types of user exploration, we propose multifaceted generative capabilities through outcome-conditioned generation. Across offline and online evaluation, we demonstrate that customizing item recommendations to the desired outcome allows for designers to provide recommendations that successfully result in those specified outcomes.
  • Windowed, Multi-Token Prediction: We propose an efficient multi-token prediction objective tailored to the structure of the sequential recommendation task, where prior next-token objectives have often lacked representational power. Instead, by optimizing multiple per-step recommendations with respect to a diverse set of future targets, we achieve notable offline and online benefits in performance, diversity, and efficiency.
@@ 2. Related Work
2.1. Sequential Recommendation Systems
Sequential recommendation systems operate on a sequence of items with which a user has interacted (Kang and McAuley, 2018a). Earlier work in academia has used recurrent neural networks (RNN) (Hidasi et al., 2016), but later work has often adopted transformers (Kang and McAuley, 2018a; Sun et al., 2019). Recent studies have harnessed the extensive knowledge of pre-trained language models (Hou et al., 2022; Li et al., 2023).
In industrial applications, a common approach is to apply sequential encoders to learn user representations (Pancha et al., 2022). The user representation can then be used in traditional recommender systems (e.g., embedding-based retrieval or ranking models). Due to its seamless integration with existing recommendation models, this approach has been widely adopted industry-wide (Hansen et al., 2020; Xu et al., 2022; Xia et al., 2023; Si et al., 2024).
Another approach in sequential recommendation systems is the generative approach, wherein the model learns to generate the items to recommend. Rajput et al. (2023) used transformers to generate items by generating a sequence of semantic ID tokens that are obtained by Residual Quantization with Variational Auto Encoder (RQ-VAE) (Lee et al., 2022; Singh et al., 2024). Zhai et al. (2024) augmented the sequence by adding heterogeneous items and also showed that the models can perform two tasks: ranking and retrieval. However, these studies have primarily concentrated on generating a single item rather than exploring the generation of sequences. In contrast, our research seeks to generate a sequence of multiple items to fully realize the potential of the generative approach. Our method is capable of recommending a diverse set of items that can drive various types of user actions. While similar approaches have been explored in academic settings (Liu et al., 2024), such strategies have not yet been widely adopted in industry.
2.2. Multi-Objective Generation
Given the multifaceted nature of user exploration and business needs on a platform like Pinterest, we explore generation techniques that consider multiple objectives. While reinforcement learning (RL) algorithms have demonstrated this capability in industry settings (Xin et al., 2020), research in language modeling indicates that RL often requires significant amounts of tuning for optimal generative capability (Ouyang et al., 2022). Recent research has introduced more streamlined and scalable techniques grounded in supervised learning. These approaches employ goal-conditioned methods (Ghosh et al., 2021) or utilize prompting strategies (Xin et al., 2022). Building on these, we develop an outcome-conditioned sequential recommendation system to serve billions of items and millions of users.
@@ 3. Problem Setup
In this section, we outline the problem setup for the task of a sequential recommendation system for a set of users $\mathcal{U}$, items $\mathcal{I}$, and actions $\mathcal{A}$. Our aim is to produce an ordered set of item recommendations $R(u,t)\subset\mathcal{I}$ of size $N$ for a user $u\in\mathcal{U}$ at a given timestep $t$. Each item $i$ has an item type $T(i)$, accounting for all kinds of items that the user can interact with (e.g., Pins). We define the set of all item types $\mathcal{T}$ as $\mathcal{T}=\{T(i)\mid i\in\mathcal{I}\}$.
Conditioned on the user $u$, we denote their historical user activity sequence ending at timestep $t_{\mathrm{max}}$ as $H(u,t_{\mathrm{max}})$, a chronologically-ordered heterogeneous sequence of items. For each item in the sequence $i_{t}\in H(u,t_{\mathrm{max}})$, the item $i_{t}$ is interacted by user $u$ at timestep $t\leq t_{\mathrm{max}}$ and its item type $T(i)\in\mathcal{T}$. We assume each interaction with an item $i_{t}$ corresponds to an action $a(i_{t})\in\mathcal{A}$.
To learn a sequential recommendation system, we propose a sequence model $f_{\theta}$ parameterized by $\theta$. To optimize the model parameters $\theta$, we leverage a set of features $\mathcal{F}(t)$ for each item type $t\in\mathcal{T}$. These may either be learned end-to-end with the sequence model or provided as learned or raw signals. The sequence model is consequently optimized using these features across the set of user sequences $H(u,t_{\mathrm{max}})$ for all $u\in\mathcal{U}$.

### v4

@@ Abstract.
This paper introduces PinRec, a novel generative retrieval model developed for applications at Pinterest. PinRec utilizes outcome-conditioned generation, enabling modelers to specify how to balance various outcome metrics, such as the number of saves and clicks, to effectively align with business goals and user exploration. Additionally, PinRec incorporates multi-token generation to enhance output diversity while optimizing generation. Our experiments demonstrate that PinRec can successfully balance performance, diversity, and efficiency, delivering a significant positive impact such as +2% sitewide clicks and +4% search repins. This paper marks a significant milestone in generative retrieval, as it presents, to our knowledge, the first rigorous study on implementing generative retrieval at the scale of Pinterest.
@@ 1. Introduction
Pinterest is a visual discovery platform where over 550 million monthly active users search for, save, and shop for ideas, represented by Pins (pins). Given the massive volume of Pins — for instance, users save over 1.5 billion Pins per week (pins2) — it is essential to provide personalized recommendations in Pinterest feeds and search results. To achieve this, recommender systems often leverage a two-stage approach: retrieval and ranking. Retrieval models select thousands of potentially relevant candidate items, then ranking models prioritize the top items to display to the user. Since ranking models depend on the retrieval results, it is important for retrieval models to identify diverse, fresh, and relevant items.
@@
Traditionally, retrieval methods employ “two-tower” models. The query tower computes query embeddings, while the item tower computes item embeddings. These models retrieve items with the highest similarity scores given a query embedding. As an alternative to two-tower models, recent work has proposed generative retrieval. Generative retrieval models leverage a sequence of tokens using generative models, such as transformers (vaswani2017attention), and then use those tokens to retrieve items. Generative retrieval has shown promising results in academic settings given its powerful ability to understand user activity history and generate candidates that existing two-tower models may not be able to retrieve (KangM18; Recformer23; UnifyingGenRetrieval24).
@@
Using generative retrieval models for large scale recommender systems poses many challenges at the scale of Pinterest. Firstly, the generative model should be capable of efficiently optimizing for different engagement outcomes such as maximizing social actions like saves and comments versus passive engagements like clicks, because business needs are constantly evolving to optimize combinations of different metrics. Secondly, the generated item sequences must possess sufficient diversity to facilitate users’ exploration of varied ideas; one-step generation of a single item representation may not suffice. Thirdly, given their increased complexity and sequential generation, generating a set of recommended items requires higher computational cost than computing an embedding with two-tower models.
@@
To our knowledge, there has been no large scale generative retrieval model that has addressed all of the aforementioned challenges. Most methods developed within academic settings (KangM18; UniS4Rec22; Recformer23) have not proven effective in meeting the latency and cost requirements necessary to handle web-scale traffic. While there are a few recent studies to apply generative retrieval to production, they focus on optimizing one objective (HSTU24; Tiger23) and are not capable of optimizing multiple target outcomes. Moreover, these methods do not address the challenge of generating a sequence of diverse items. Because generating multiple items sequentially is computationally expensive, existing approaches typically generate only the next item rather than a full set at once.
Figure 2. Modeling diagram of PinRec
@@
In this paper, we present PinRec, a generative retrieval technique for real-world recommender systems (Figure 2). PinRec builds on other generative retrieval models, leveraging transformers to perform next token prediction (KangM18; HSTU24; Tiger23). Unlike previous methods, however, PinRec addresses and resolves three aforementioned challenges for generation of optimal recommendations: multi-objective generative capabilities, computational efficiency, and item diversity. Our main contributions in the paper are summarized below.
  • Outcome-Conditioned Generative Retrieval: PinRec employs outcome-conditioned generation, enabling the optimization of multiple business metrics. For each outcome (e.g., business metrics such as user saves or clicks), PinRec generates the next item token that is most likely to maximize that specific outcome, and then retrieves candidate items to balance these objectives according to the desired trade-offs (Fig. 2). Through both offline and online evaluations, we show that tailoring item recommendations to particular outcomes enables designers to effectively achieve those specified objectives.
  • Multi-Token Generation for Efficient Sequence Generation: One obstacle to generating an entire sequence of items for retrieval is that traditional models typically produce only one item at a time. To address this, we propose a multi-token generation framework (Multitoken24), in which PinRec generates multiple future items in parallel, as illustrated in Figure 2. This approach significantly reduces the number of steps needed to produce diverse output embeddings, resulting in substantial offline and online improvements in performance, diversity, and efficiency.
@@
We will make the model code publicly available upon publication.
@@ 2. Related Work
Sequential Recommendation Systems: Sequential recommendation systems operate on a sequences of items with which a user has interacted (KangM18). Earlier work in academia has used recurrent neural networks (RNN) (GRU4REC16), but later work has often adopted transformers (KangM18; BERT4REC19). Recent studies have harnessed the extensive knowledge of pre-trained language models (UniS4Rec22; Recformer23).
In industrial applications, a common approach is to apply sequential encoders to learn user representations (Pinnerformer22). The user representation can then be used in traditional recommender systems (e.g., embedding-based retrieval or ranking models) (Spotify20; rethink; Transact23; TWINV224). Another approach in sequential recommendation systems is the generative approach, where the model learns to generate the items to recommend. Tiger23 used transformers to generate items by producing a sequence of semantic ID tokens (RQVAE22; SemanticIDGeneral24). HSTU24 also proposed a generative retrieval method. However, these studies have mainly focused on generating a single item to optimize a single action. In contrast, our research aims to generate sequences of multiple items to optimize for multiple outcomes.
Multi-Objective Generation: We discuss generative models that consider multiple objectives. While RL algorithms have shown this capability in industry (SelfsupervisedRL20), studies in language modeling suggest RL often needs extensive tuning for optimal generation (PPO22). Recent research has introduced more streamlined and scalable techniques based on supervised learning, such as goal-conditioned methods (IterativeSupervised21) and prompting strategies (PromptingRL22). PinRec builds on these approaches to enhance scalability.
@@ 3. Problem Setup
We formulate the generative sequential recommendation task as follows. Consider a system with users $\mathcal{U}$, items $\mathcal{I}$, actions $\mathcal{A}$, and surfaces $\mathcal{S}$. Our objective is to generate an ordered list of $N$ item recommendations $R(u,t)\subset\mathcal{I}$ for user $u\in\mathcal{U}$ at timestep $t$.
Item Types and Actions. Each item $i\in\mathcal{I}$ belongs to an item type $T(i)\in\mathcal{T}$, where $\mathcal{T}=\{T(i)\mid i\in\mathcal{I}\}$ represents the set of all item types in the system (e.g., Pins, boards, search queries). Users can perform various actions $\mathcal{A}=\{\text{click},\text{save},\text{share},\text{comment},\ldots\}$ on these items across different surfaces. Each interaction occurs on a specific surface $s\in\mathcal{S}$, where surfaces represent different recommendation products, such as homefeed, search results, board pages, and related pin feed.
User History. For each user $u$, we maintain their interaction history $H(u,t_{\text{max}})=\langle(i_{1},a_{1},s_{1},t_{1}),(i_{2},a_{2},s_{2},t_{2}),\ldots,(i_{m},a_{m},s_{m},t_{m})\rangle$, where each tuple $(i_{j},a_{j},s_{j},t_{j})$ represents an interaction with item $i_{j}\in\mathcal{I}$ via action $a_{j}\in\mathcal{A}$ on surface $s_{j}\in\mathcal{S}$ at timestep $t_{j}\leq t_{\text{max}}$. The sequence is chronologically ordered such that $t_{1}\leq t_{2}\leq\ldots\leq t_{m}$.
Sequential Recommendation Model. We propose a parameterized sequence model $f_{\theta}:\mathcal{H}\times\mathcal{T}\times\mathcal{S}\rightarrow\mathbb{R}^{|\mathcal{I}|}$ that maps a user’s interaction history, item types, and surface context to item scores, where $\mathcal{H}$ denotes the space of all possible interaction histories. The model parameters $\theta$ are optimized over the collection of user interaction sequences $\{H(u,t_{\text{max}})\mid u\in\mathcal{U}\}$ to maximize the likelihood of observed user interactions while accounting for the heterogeneous nature of actions and surfaces. This enables the model to learn nuanced behavioral patterns that depend on both what users interact with and how and where they interact with it.